1,320 crore s2 Calculate (a) Net National Product at Market Price and (b) Private 100 from the following: (in crores) 800 30 600 Items Net current transfers to abroad 250 40 ) Mixed income (m) Subsidies iv) Operating surplus )National debt interest )Net factor income to abroad 20 130 200 70 10 1400 vi) Compensation of employees Indirect tax ) Domestic product accruing to government Current transfers by government. 100 350 50
Answer #1

a. NNPmp = compensation of enployees+ Operating surplus+mixed income+(Indirect tax - subsidies)-net factor income to abroad.

=1400+200+600(100-20) - 10

=2200+80-10

=₹2270 crores

b. Private income= NNPmp- net indirect tax- domestic product accruing to government- net current transfers to abroad+national debt interest+current transfers by government.

=2270-(100-20)-350-30+70+50

=₹1930 crores